OpenStudy (anonymous):

For the equation given below, evaluate dy/dx at the point (-2,2). 2y^3+y^2-6x^2 = -4

OpenStudy (anonymous):

diff. w.r.t x \[6y ^{2}\frac{ dy }{dx }+2y \frac{ dy }{dx }-12x=0\] \[y \left( 6y+2 \right)\frac{ dy }{dx }=12x\] put x=-2,y=2 and find the value of \[\frac{ dy }{ dx }\]